OpenStudy (mimi_x3):

Two sections are cut out from a wire of length 25cm. The sections are then used to form two different circles. Find the lowest total area of the circles.

OpenStudy (anonymous):

x + y = 25 where x and y are the circumferences of 2 circles A =[( x /2pi)^2]*pi + [( y /2pi)^2]*pi A = x^2/4pi + y^2/4pi now from first equation y = 25-x plug this into the above equation then differentiate and equate to 0 solve for x and find y calculate the areas using these values

hero (hero):

Good job Jimmy

OpenStudy (anonymous):

thanx!

OpenStudy (mimi_x3):

Thank you (: